If I drive x mph for t hours and then y mph for 2t hours, what is the expression for average speed?
Answer by stanbon(75887)

If I drive x mph for t hours and then y mph for 2t hours, what is the expression for average speed?
------------------
average speed = (total distance)/(total time)
distance = rate*time
----------------------
Partial distance: xt miles
Partial distance: y*2t miles
----------------------------------
Total time = t + 2t
------------------------------------
Equation:
Average speed = (xt+2yt)/3t = (x+2y)/3 mph
=================================================
Cheers,
Stan H.